# LM-MCVT: A Lightweight Multi-modal Multi-view Convolutional-Vision Transformer Approach for 3D Object Recognition

### v2

@@ I Introduction
In general, 3D object recognition approaches can be categorized into three main categories: voxel-based, point-based, and view-based methods. Voxel-based methods [2, 3, 4] and point-based methods [5, 6, 7, 8] use 3D meshes and point clouds, respectively. While these approaches comprehensively capture the spatial information of 3D objects, their high computational costs limit practicality [9]. In contrast, view-based methods enhance efficiency by converting 3D objects into 2D images from various viewpoints, simplifying the process and eliminating complex 3D feature processing [10]. For example, as shown in Fig. 1, a mug with rich features can be accurately recognized from multiple viewpoints. Benefiting from the advancement of deep learning, various methods encoded the rendered views to achieve a robust representation of a given object. For instance,  [11, 12, 13] utilize 2D CNNs to extract features from multiple views of an object and aggregate these features for 3D object classification purposes. However, CNNs primarily focus on capturing local details while often overlooking global contextual relationships. Inspired by the ability of vision transformers to capture global spatial information,  [14] introduced the multi-view vision transformer (MVT) for enhanced 3D object recognition. However, MVT relies on a simple average-pooling operation for view aggregation, which overlooks variations and unique contributions among views. Furthermore, as the number of ViT encoders increases, the model’s ability to retain local features weakens, limiting its overall performance. Previous research on multi-view object recognition was mainly focused on dense multi-view techniques to achieve better accuracy. While increasing the number of views can to some extend improve recognition accuracy, acquiring dense multi-view data in real-world applications is often constrained by computational and memory requirements. Additionally, single-view methods incorporating depth information have been introduced to enhance object representation for 3D object recognition [15]. Compared to dense multi-view approaches, these single-view methods offer better efficiency but still face limitations in stability and robustness [15].
@@
  • We develop the Globally Entropy-based Embeddings Fusion (GEEF) method, a novel feature aggregation strategy that leverages entropy-based weighting to adaptively fuse view-specific embeddings, which efficiently integrates multi-modal multi-view representations for 3D object recognition.
@@ II Related work
View-based methods: These methods represent a 3D object by projecting it into multiple 2D images captured from various viewpoints. Our method belongs to this category. Hou et al. [17] first used CNNs to encode multiple views, then applied reinforcement learning to select the next-best-view for 3D object recognition. Alzahrani et al. [12] used the most discriminative views of an object and achieved competitive accuracy in 3D object classification tasks. While these approaches have advanced the field, their reliance on heavy backbone networks often results in high memory usage and increased computational time, posing challenges for mobile devices. Moreover, such methods frequently prioritize extracting local features over capturing global spatial information [18]. Chen et al. [14] proposed the Multi-view Vision Transformer (MVT), a 3D object classification network that uses the middle-weight DeiT backbone. MVT incorporates multi-head attention to capture global object information and enables effective communication across different views through a unified global structure. However, as the number of layers in the MVT backbone increases, the focus on capturing local details from input views diminishes [18]. To further enhance 3D object recognition, Wang et al. proposed the optimal viewpoint pooling transformer, which obtains the best viewpoint settings from dense views to optimize performance [13]. In contrast to our approach, this method faces substantial challenges when optimal viewpoints are unavailable. Obtaining all views or dense views for a mobile robot in real-world scenarios, particularly in dynamic environments, presents practical difficulties.
@@ III-E Globally Entropy-based Embeddings Fusion (GEEF)
Next, the entropy of the RGB representation, denoted as $H^{\text{cls}}_{v}(c_{v_{j}})=-\sum_{j=1}^{l}p_{c_{v_{j}}}\log p_{c_{v_{j}}}$, is computed from the class token vectors across all RGB views, aggregated in s $\mathbf{C}_{v}^{rgb}=[\mathbf{c}_{v_{1}},\mathbf{c}_{v_{2}},\ldots,\mathbf{c}_{v_{j}},\ldots,\mathbf{c}_{v_{l}}]$. To utilize entropy values for feature fusion, they need to be normalized to obtain the weights $\mathbf{w}_{v_{j}}$ for each view. The normalization formula is as follows:
@@ IV-A Implementation Details
We used two datasets: the synthetic ModelNet dataset [22] and the real-world OmniObject3D dataset [23]. For the offline experiments on ModelNet, which includes ModelNet10 and ModelNet40, we evaluated our model using the training and testing splits outlined in [14]. ModelNet40 comprises $12311$ 3D CAD models divided into 40 categories, with $9843$ models for training and $2468$ for testing. ModelNet10, a subset, consists of 10 categories. OmniObject3D is a large-scale dataset of high-quality, real-scanned 3D objects, featuring $6000$ objects across $190$ everyday categories. For OmniObject3D, we adopted the same view settings as ModelNet. To evaluate our approach, we employed a 5-fold cross-validation strategy, which provides a robust estimate of the model’s performance and minimizes the risk of overfitting to specific subsets. We conducted our experiments using an NVIDIA V100 graphics card. Our approach was trained with the Adam optimizer, starting with a learning rate of $0.0002$ and betas set to [$0.9,0.98$]. To optimize training, we dynamically adjusted the learning rate using a cosine annealing schedule.
@@ IV-C Ablation Study for Convolutional Encoders and Global Transformers
TABLE II: Ablation Study on the Impact of Convolutional Encoders and Transformers on Model Accuracy (%) with Multi-Modal (RGBD) Data
@@ IV-E1 Performance on ModelNet Dataset
As summarized in Table VI, voxel-based methods, such as 3DShapeNet and VoxNet, show limited performance, achieving a maximum accuracy of 92.0% on ModelNet10 and 83.0% on ModelNet40. These approaches struggle with computational efficiency and resolution loss due to the voxelization process, which limits their ability to capture fine-grained details of 3D objects. Point-based methods like PointTramba and PCNN outperform voxel-based methods by directly operating on point clouds, achieving up to 94.9% accuracy on ModelNet10. However, these methods still fall short compared to view-based approaches due to challenges in efficiently capturing global and contextual features from sparse point clouds. View-based methods consistently outperform voxel-based and point-based techniques, leveraging multiple 2D projections to better preserve object details and contextual relationships. Among these, methods like MVCNN and MVT achieve high accuracy but at the cost of significantly increased computational requirements due to the use of dense multi-view setups. Our method, LM-MCVT, achieves the best performance while maintaining computational efficiency. Specifically, on ModelNet10, LM-MCVT achieves 98.5% accuracy using only four views, outperforming heavier methods such as MVCNN, which requires 80 views, and lightweight approaches such as MVT and TMTL, which use 12 views. On ModelNet40, our method reaches an accuracy of 95.6%, demonstrating its robustness across datasets. Our method sets a new benchmark in terms of accuracy and efficiency. In particular, by effectively combining RGB and depth information, LM-MCVT leverages complementary features from both modalities, providing a richer representation of 3D objects. Furthermore, despite using only 10.5M parameters and four views, our method achieves state-of-the-art performance, offering a practical solution for real-world applications where computational resources and viewpoints are often limited. We attributed these results to the point that LM-MCVT extracted both local and global features through its convolutional and transformer-based architecture, allowing it to outperform methods with significantly more parameters or views.
@@
IV-E2 Performance on OmniObject3D Dataset [23]

### v3

@@ I Introduction
In general, 3D object recognition approaches can be categorized into three main categories: voxel-based, point-based, and view-based methods. Voxel-based methods [2, 3, 4] and point-based methods [5, 6, 7, 8] use 3D meshes and point clouds, respectively. While these approaches comprehensively capture the spatial information of 3D objects, their high computational costs limit practicality [9]. In contrast, view-based methods enhance efficiency by converting 3D objects into 2D images from various viewpoints, simplifying the process and eliminating complex 3D feature processing [10]. For example, as shown in Fig. 1, a mug with rich features can be accurately recognized from multiple viewpoints. Benefiting from the advancement of deep learning, various methods encoded the rendered views to achieve a robust representation of a given object. For instance,  [11, 12, 13] utilize 2D CNNs to extract features from multiple views of an object and aggregate these features for 3D object classification purposes. However, CNNs primarily focus on capturing local details while often overlooking global contextual relationships. Inspired by the ability of vision transformers to capture global spatial information,  [14] introduced the multi-view vision transformer (MVT) for enhanced 3D object recognition. However, MVT relies on a simple average-pooling operation for view aggregation, which overlooks variations and unique contributions among views. Furthermore, as the number of ViT encoders increases, the model’s ability to retain local features weakens, limiting its overall performance. Previous research on multi-view object recognition was mainly focused on dense multi-view techniques to achieve better accuracy. Increasing the number of views can, to some extent, improve recognition accuracy, but acquiring dense multi-view data in real-world applications is often constrained by computational and memory requirements. Additionally, single-view methods incorporating depth information have been introduced to enhance object representation for 3D object recognition [15]. Compared to dense multi-view approaches, these single-view methods offer better efficiency but still face limitations in stability and robustness [15].
@@
  • We develop the Globally Entropy-based Embeddings Fusion (GEEF) method, a novel feature aggregation strategy that leverages entropy-based weighting to adaptively fuse embeddings from multiple viewpoints. This approach efficiently integrates multi-modal multi-view representations, enhancing the performance of 3D object recognition.
@@ II Related work
View-based methods: These methods represent a 3D object by projecting it into multiple 2D images captured from various viewpoints. Our method belongs to this category. Hou et al. [17] used CNNs to encode multiple views, then applied reinforcement learning to select the next-best-view for 3D object recognition. Alzahrani et al. [12] employed the most discriminative views of an object and achieved competitive accuracy in 3D object classification tasks. While these approaches have advanced the field, their reliance on heavy backbone networks often results in high memory usage and increased computational time, posing challenges for mobile devices. Moreover, such methods frequently prioritize extracting local features over capturing global spatial information [18]. Chen et al. [14] proposed the Multi-view Vision Transformer (MVT), a 3D object classification network that uses the middle-weight DeiT backbone. MVT incorporates multi-head attention to capture global object information and enables effective communication across different views through a unified global structure. However, as the number of layers in the MVT backbone increases, the focus on capturing local details from input views diminishes [18]. To further enhance 3D object recognition, Wang et al. proposed the optimal viewpoint pooling transformer, which obtains the best viewpoint settings from dense views to optimize performance [13]. In contrast to our approach, this method faces substantial challenges when optimal viewpoints are unavailable. Obtaining all views or dense views for a mobile robot in real-world scenarios, particularly in dynamic environments, presents practical difficulties.
@@ III-E Globally Entropy-based Embeddings Fusion (GEEF)
Next, the entropy of the RGB representation, denoted as $H^{\text{cls}}_{v}(c_{v_{j}})=-\sum_{j=1}^{l}p_{c_{v_{j}}}\log p_{c_{v_{j}}}$, is computed from the class token vectors across all RGB views, aggregated in $\mathbf{C}_{v}^{rgb}=[\mathbf{c}_{v_{1}},\mathbf{c}_{v_{2}},\ldots,\mathbf{c}_{v_{j}},\ldots,\mathbf{c}_{v_{l}}]$. To utilize entropy values for feature fusion, they need to be normalized to obtain the weights $\mathbf{w}_{v_{j}}$ for each view. The normalization formula is as follows:
@@ IV-A Implementation Details
We used two datasets: the synthetic ModelNet dataset [22] and the real-world OmniObject3D dataset [23]. For the offline experiments on ModelNet, which includes ModelNet10 and ModelNet40, we evaluated our model using the training and testing splits outlined in [14]. ModelNet40 contains $12311$ 3D CAD models across 40 categories, with $9843$ models for training and $2468$ for testing. ModelNet10 is a 10-category subset. OmniObject3D is a large-scale dataset of high-quality, real-scanned 3D objects, featuring $6000$ objects across $190$ everyday categories. For OmniObject3D, we adopted the same view settings as ModelNet. To evaluate our approach, we employed a 5-fold cross-validation strategy, which provides a robust estimate of the model’s performance and minimizes the risk of overfitting to specific subsets. We conducted our experiments using an NVIDIA V100 graphics card. Our approach was trained with the Adam optimizer, starting with a learning rate of $0.0002$ and betas set to [$0.9,0.98$]. To optimize training, we dynamically adjusted the learning rate using a cosine annealing schedule.
@@ IV-C Ablation Study for Convolutional Encoders and Global Transformers
TABLE II: Ablation Study on the Impact of Convolutional Encoders and Transformers on Model Accuracy (%)
with Multi-Modal (RGB-D) Data
@@ IV-E1 Performance on ModelNet Dataset
As summarized in Table VI, voxel-based methods, such as 3DShapeNet and VoxNet, show limited performance, achieving a maximum accuracy of 92.0% on ModelNet10 and 83.0% on ModelNet40. These approaches struggle with computational efficiency and resolution loss due to the voxelization process, which limits their ability to capture fine-grained details of 3D objects. Point-based methods like PoinTramba and PCNN outperform voxel-based methods by directly operating on point clouds, achieving up to 94.9% accuracy on ModelNet10. However, these methods still fall short compared to view-based approaches due to challenges in efficiently capturing global and contextual features from sparse point clouds. View-based methods consistently outperform voxel-based and point-based techniques, leveraging multiple 2D projections to better preserve object details and contextual relationships. Among these, MVCNN achieves good accuracy but at the cost of significantly increased computational requirements due to the use of dense multi-view setups. Our method, LM-MCVT, achieves the best performance while maintaining computational efficiency. Specifically, on ModelNet10, LM-MCVT achieves 98.5% accuracy using only four views, outperforming heavier methods such as MVCNN, which requires 80 views, and lightweight approaches such as MVT and TMTL, which use 12 views. On ModelNet40, our method reaches an accuracy of 95.6%, demonstrating its robustness across datasets. Our method sets a new benchmark in terms of accuracy and efficiency. In particular, by effectively combining RGB and depth information, LM-MCVT leverages complementary features from both modalities, providing a richer representation of 3D objects. Furthermore, despite using only 10.5M parameters and four views, our method achieves state-of-the-art performance, offering a practical solution for real-world applications where computational resources and viewpoints are often limited. We attributed these results to the point that LM-MCVT extracted both local and global features through its convolutional and transformer-based architecture, allowing it to outperform methods with significantly more parameters or views.
@@
IV-E2 Performance on OmniObject3D Dataset